OpenStudy (anonymous):

FAN AND MEDAL i really need help especially with #3 1. Write the inequality for the graph? A. y ≥ –2x + 4 B. y > –2x + 4 C. y < –2x + 4*MY CHOICE* D. y ≤ –2x + 4 2. What are the x- and y-intercepts of the line: 2x – 3y = –12? A. x-intercept –4, y-intercept 6 B. x-intercept 4, y-intercept –6 C. x-intercept 6, y-intercept –4 D. x-intercept –6, y-intercept 4 3.Write the equation of the graph? A. y – 3 = ½(x + 1) B. y + 3 = –½(x – 1) C. y – 3 = –½(x + 1) D. y = ½x + 5

OpenStudy (anonymous):

I thought A or B For the first one..

OpenStudy (anonymous):

@ClarissaMorgenstern hmmm... i dont think so because a dashed boundary line with shading below the boundary line uses the < symbol but i'm not sure

OpenStudy (anonymous):

For the first one, your choice is correct. Plug in the point (0,0) into the inequality- you'll see it works.

OpenStudy (anonymous):

For the second one, first rewrite the equation to isolate y. 2x – 3y = –12 -3y = -12 -2x y = 4 +2/3x

OpenStudy (anonymous):

Setting y to 0 gives us: 0 = 4 + 2/3x x = 6

OpenStudy (anonymous):

Setting x to 0 gives us : y = 4 + 2/3(0) y= 4

OpenStudy (anonymous):

@greenglasses thank you

OpenStudy (anonymous):

For the last one, the answer is D- it's the only one with a y-intercept of 5.

OpenStudy (anonymous):

@greenglasses thanks i really do appreciate it

OpenStudy (anonymous):

Np